OpenStudy (anonymous):

If a person walks with a speed of 1.3 miles per hour, it takes 135 minutes to walk around the rim. Assuming that the rim is approximately a circle, what is the diameter of the crater? (Round your answer to the nearest tenth and make sure you convert the minutes to hours first!)

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (shinalcantara):

use the formula for circumference and convert minutes to hours for you to be able to get the distance travelled

OpenStudy (anonymous):

I did that and it said the answer is wrong

OpenStudy (cwrw238):

135 minutes = 2 hours 15 minutes = 2.25 hours speed = distance / time so distance walked= speed * time = 2.25* 1.3 = 2.925 miles

OpenStudy (cwrw238):

diameter of the crater = 2.925 / pi = 0.93 or 0.9 mile to nearest tenth

OpenStudy (anonymous):

Thank you!

OpenStudy (cwrw238):

yw